# Cluster Synchronization via Graph Laplacian Eigenvectors

### v1

@@ III Stability of Structurally motivated Clustering
In the previous sections we have shown how with the coefficient equation, we can frame cluster synchronization as the dynamics of AEP structural eigenvectors of the graph Laplacian. The natural question is then what other kinds of clustering the Laplacian basis could allow for. Framing this problem in the linear algebra of an eigenbasis and AEP provides a natural direction for loosening the structural restrictions for synchronized clusters. Strict AEPs are rare and the associated full phase synchronization within clusters may be stronger than what is required for application. The concept of a weakened AEP has been considered previously, such as in the model of stochastically equitable partitions [10], and even in the setting of the Kuramoto model with a bound on the out degrees between partitions [22]. We approach this question from the perspective of the spectral theory behind an AEP to see how sensitive the eigenbasis is to perturbation, and how a spectral theory around its weakening could fit into the spectral description of synchronization presented in this paper.
@@ III.1 Approximate Almost Equitable Partitions
In theorem 1 we saw that given a partition $\pi$ with indicator matrix $P$ on a graph with Laplacian $L$, a sufficient and necessary condition for $\pi$ to be an AEP is for $LP=PL^{\pi}$. The entries of $LP$ correspond to the out edge weight sums of a vertex to each cluster of $\pi$. Meanwhile, the matrix $PL^{\pi}$ inhabits the same form, but assigns to each vertex the average weight sum of its cluster to each other cluster. Thus we define $E=PL^{\pi}-LP$ which captures the deviation of a vertex weight sums from the averages and allows us to approximate the spectral AEP property. The minimization of a similar matrix was explored in [26] in the context of finding regular equitable partitions.
@@
$E$ $=L^{\prime}P-(L^{\prime})^{\pi}P$
$=(L+N)P-P(P^{T}P)^{-1}P^{T}(L+N)P$
$=LP-P(P^{T}P)^{-1}P^{T}LP+NP-P(P^{T}P)^{-1}P^{T}NP$
$=NP-P(P^{T}P)^{-1}P^{T}NP$
$=NP-N^{\pi}P.$
When put in this form, the Laplacian L, which supports an AEP cancels out. This shows how $E$ can be described in terms of the deviation from a graph Laplacian supporting the AEP. Thus it is the placement of the noise over the graph that determines the equitable error. This suggests that the probability distribution over which error is introduced could have profound impact on how it is represented in the equitable error (Appendix C explores one such avenue). We can bound the norm of $\epsilon$ and relate the values of $E$ to the size of $\epsilon$ more explicitly.
@@
Let $G$ be a weighted graph with graph Laplacian $L$. Let $\pi$ be a vertex partition. The equitable error of any eigenvector $v$ of $L^{\pi}$ is bounded as $\|\epsilon\|\leq\sigma_{1}(E)\|v\|\leq 2k\|v\|\max_{i}\sum_{j}\left|e_{ij}\right|$
@@
Recall that for a rectangular matrix $A$ the largest singular value, $\sigma$, bounds its products, $\|Av\|\leq\sigma\|v\|$. It can be shown that $\max_{x\not=0}\frac{\|Ax\|_{\infty}}{\|x\|_{\infty}}=\max_{i}\sum_{j}|a_{ij}|$. The largest singular value of a matrix is equal to its operator norm. Thus it follows that
@@
$\|\epsilon\|\leq\sigma_{1}(E)\|v\|\leq 2k\|v\|\max_{i}\sum_{j}\left|e_{ij}\right|$
The size of this bound is determined by the vertex which differs the most in magnitude from the average in out-degrees. It is also of note that this bound scales with the number of clusters in the partition, $k$. Thus fewer partitions de-emphasize the maximum deviations from the average. This bound represents the absolute worst case scenario, and is not tight. In reality, the vectors $v$ have no reason to be exactly in line with the largest singular vector, so we can expect better results in application.
@@
Let $G$ be a weighted, undirected graph with graph Laplacian $L$. Let $\pi$ be a vertex partition such that the equitable error $\|\epsilon\|\leq\delta$ for eigenvector-value pair $(\lambda,v)$ of $L^{\pi}$. Then $Pv$ is the best approximated by eigenvectors with eigenvalues close to $\lambda$. For any truncated approximation, $u$ using eigenvectors with eigenvalues within $\gamma$ of $\lambda$, the bound $\|Pv-u\|\leq(\delta/\gamma)\sqrt{n-m}$ will hold.
@@
Consider the basis of eigenvectors of $L$. The Laplacian is symmetric and therefore diagonalizable by the undirectedness of the graph. Therefore its eigenvalues form an orthogonal basis. Let $\{v_{1},\dotsc,v_{n}\}$ be a set of normal representatives, forming the spectral basis of $L$.
@@
[…]
Note that $L$ is symmetric and diagonally dominant, and therefore has non negative eigenvalues. From the above result we have that $\beta_{i}=(\lambda-\lambda_{i})\alpha_{i}$ for $1\leq i\leq n$. Thus $|(\lambda-\lambda_{i})\alpha_{i}|\leq\delta$ for all $i$.
It follows that when $\delta$ is small, $(\lambda-\lambda_{i})$ or $\alpha_{i}$ is small. For a fixed $\beta_{i}$ value, the further away $\lambda_{i}$ is from $\lambda$, the smaller $\alpha_{i}$ must be. This means that eigenvectors of $L$ with eigenvalues close to $\lambda$ can have larger coefficients $\alpha_{i}$ and therefore contain the most information of $w$. Let $A=\{a_{1},\cdots,a_{m}\}$ be the set of coefficients for eigenvalues within $\gamma$ of $\lambda$. For all other eigenvalues, $\alpha_{i}$ is likely small. Therefore $u=\alpha_{a_{1}}v_{a_{1}}+\cdots+\alpha_{a_{m}}v_{a_{m}}$ is our description of the eigenvector ‘close’ to $Pv$. This allows for the derivation of a formal bound.
@@
This characterizes an eigenvector whose eigenvalue is close to that of $Pv$ as being a good approximation of it. As $\|\epsilon\|$ approaches zero, $\|Pv-u\|$ does as well for any truncated approximation $u$. This implies a stability in the spectral properties of the almost equitable partition. We call these approximations of AEPs, Quasi Equitable Partitions (QEPs). Whereas [22] define their weakening in terms of a bound on the out degree of vertices, we define ours as a deviation from the average, which allows it to maintain a relationship with the underlying spectral theory as we have demonstrated above.
@@ III.2 Weak Cluster Synchronization
Because our description of cluster synchronization is spectral instead of combinatorial, we can immediately apply the same methodology from AEPs to the structural eigenvectors of QEPs. When moving from an AEP to a QEP, we can no longer construct the full eigenbasis of a corresponding quotient Laplacian. Non-structural eigenvectors have nonzero coefficients in the eigendecomposition of $\vec{\omega}$. Information is leaked into the rest of the eigenbasis, and it is precisely this leakage which results in differences in phase within clusters. This is small if the QEP is reasonably close to an AEP. In this case the contribution of the non structural eigenvectors in the description of the natural frequencies results in a small but nonzero end state coefficient for them. Additionally, the structural eigenvectors are no longer completely constant within clusters. The non exact cluster synchronization arises from the difference in scaling moving from the decomposition of the natural frequency vector to the decomposition of the end state of the system.
Figure 5: The graph A. depicts a weighted network of $15$ vertices, and a color coded QEP which it admits. A phase plot is shown in B. indicating the phase locked cluster synchronization of the Kuramoto model placed on this graph. Note the slight spread of phases within clusters. Finally, observe in C. the cluster synchronized state is described almost entirely by the quotient graph sub-basis, $\alpha_{1}(t)$ and $\alpha_{2}(t)$, with some leakage of information to the rest of the eigenbasis.
In section II.1.1 we discussed assortativity and spectrum placement with AEPs. The same argument using random walks can be extended to weighted graphs and QEPs [20]. The proximity of a QEP to a full AEP determines the bound of the variance in the intra cluster values of a structural eigenvector. With a reasonably small bound, we can expect generically few zero crossings to occur within a cluster.
@@ IV.1 Clustered Dynamics at Separate Frequencies
An analytically tractable way to explore this through the coefficient equations is by considering the single unstable mode dynamics. Consider a coefficient $11\leq r_{1}\leq n-$ 1, and assume non zero modes $r\neq 0,r_{1}$ have settled into the asymptotic state predicted in the linearized case $\alpha_{r}^{\infty}$. We extend the derivation from [21] to weighted graphs. One can find parameter values of a regime change for the dynamical mode $\alpha_{r_{1}}$ in a second order approximation. Particularly we see the following, the derivation of which is in appendix B.
@@
This shows that the lower spatial frequencies can support dynamical modes better. In the context of our analysis of structural eigenvectors, this suggests that the eigenvectors of assortative AEPs or QEPs allow for unstable mode behavior more easily. Thus clusters with strong intra cluster connectivity and weak inter cluster connectivity should be able to maintain separate cluster frequencies if the disparity in connectivity is high enough and the natural frequencies align well enough with one of the structural eigenvectors. This intuition also suggests that an analogous property should hold for multiple unstable modes, if they all share reasonably high alignment with the natural frequency vector.
@@
Figure 6: A. plots $\Delta_{r}$ for all the modes, $r$, of a sample graph, as functions of $\sigma$, with fixed natural frequencies. The phase plots in B. show the system experiencing multi-frequency clustered behavior. Finally, C. depicts the coefficients of the phase decomposition, indicating a single dynamical mode, $\alpha_{1}(t)$. This behavior is shown possible by the fact the corresponding discriminant in A. is negative.
This reasoning does not extend easily to QEPs. Certainly the machinery works for any eigenvectors of the graph Laplacian, but the information leak of the structural eigenvectors for a QEP complicates the description of clustered behavior. Given a system with constant natural frequencies within clusters, the structural eigenvectors of an associated QEP generically won’t be able to fully describe the clustered dynamics. In the frequency locked synchronization regime, this resulted in small activation from other modes. When considering a single unstable mode, that information from non structural modes can allow them to also become unstable. This behavior distances the dynamics from the above analysis, and makes the results less interpretable. However, a more catered choice of natural frequencies should allow for similar behavior.

### v2

@@ III Stability of Structurally motivated Clustering
In section II we have shown how with the coefficient equation, we can frame cluster synchronization in terms of the dynamics of AEP structural eigenvectors of the graph Laplacian. The natural question is then what other kinds of clustering the Laplacian basis could allow for. Framing this problem in the linear algebra of an eigenbasis and the AEP provides a natural direction for loosening the structural restrictions for synchronized clusters. Strict AEPs are rare and the associated full phase synchronization within clusters may be stronger than what is required for application. The concept of a weakened AEP has been considered previously, such as in the model of stochastically equitable partitions,Schaub, Li, and Peel (2023) and even in the setting of the Kuramoto model with a bound on the edge weight sums between partitions.Kato and Ishii (2023) We approach this question from the perspective of the spectral theory behind an AEP to see how sensitive the eigenbasis is to perturbation, and how a spectral theory around its weakening could fit into the spectral description of synchronization presented in this paper.
@@ III.1 Approximate Almost Equitable Partitions
In theorem 1 we saw that given a partition $\pi$ with indicator matrix $P$ on a graph with Laplacian $L$, a sufficient and necessary condition for $\pi$ to be an AEP is that $LP=PL^{\pi}$. The entry $(LP)_{ip}$ corresponds to the edge weight sum of vertex $i$ to all its neighbors in cluster $p$ of $\pi$. Meanwhile the matrix $PL^{\pi}$ inhabits the same form, but in the $p$th column, assigns to each vertex in a given cluster the average edge weight sum from that cluster to cluster $p$. Thus we define $E=PL^{\pi}-LP$ which captures the deviation of the vertex edge weight sums from the averages and allows us to approximate the spectral AEP property. The minimization of a similar matrix was explored inScholkemper and Schaub (2023) in the context of finding regular equitable partitions.
@@
$E$ $=L^{\prime}P-P(L^{\prime})^{\pi}$
$=(L+N)P-P(P^{T}P)^{-1}P^{T}(L+N)$
$=LP-P(P^{T}P)^{-1}P^{T}L+NP-P(P^{T}P)^{-1}P^{T}N$
$=NP-P(P^{T}P)^{-1}P^{T}N$
$=NP-PN^{\pi}.$
When put in this form, the Laplacian L, which supports an AEP cancels out. This shows how $E$ can be described in terms of the deviation from a graph Laplacian supporting the AEP. Thus it is the placement of the noise over the graph that determines the equitable error. This suggests that the probability distribution over which error is introduced could have profound impact on how it is represented in the equitable error (Appendix D explores one such avenue). We can bound the norm of $\epsilon$ and relate the values of $E$ to the size of $\epsilon$ more explicitly.
@@
Let $G$ be a weighted graph with graph Laplacian $L$. Let $\pi$ be a vertex partition. The equitable error of any eigenvector $v$ of $L^{\pi}$ is bounded as $\|\epsilon\|\leq\sigma_{1}(E)\|v\|\leq 2k\|v\|\max_{i}\sum_{j}\left|e_{ij}\right|$ where $\sigma_{1}(E)$ is the largest singular value of $E$ and $k$ is the number of cells in $\pi$.
@@
Recall that for a rectangular matrix $A$ the largest singular value, $\sigma_{1}(A)$, bounds its products, $\|Av\|\leq\sigma_{1}(A)\|v\|$. It can be shown that $\max_{x\not=0}\frac{\|Ax\|_{\infty}}{\|x\|_{\infty}}=\max_{i}\sum_{j}|a_{ij}|$. The largest singular value of a matrix is equal to its operator norm. Thus it follows that
@@
$\|\epsilon\|\leq\sigma_{1}(E)\|v\|\leq 2k\|v\|\max_{i}\sum_{j}\left|e_{ij}\right|.$
The size of this bound is determined by the vertex which differs the most in magnitude from the average out edge weight sums of its cluster. It is also of note that this bound scales with the number of clusters in the partition, $k$. Thus fewer partitions de-emphasize the maximum deviations from the average. This bound represents the absolute worst case scenario, and is not tight. In reality, the vectors $v$ have no reason to be exactly in line with the largest singular vector, so we can expect better results in application.
@@
Let $G$ be a weighted graph with graph Laplacian $L$. Let $\pi$ be a vertex partition such that the equitable error $\|\epsilon\|\leq\delta$ for eigenvector-value pair $(\lambda,v)$ of $L^{\pi}$. Then $Pv$ is best approximated by eigenvectors with eigenvalues close to $\lambda$. For any truncated approximation, $u$ using eigenvectors with eigenvalues within $\gamma$ of $\lambda$ and assuming $m$ of these, the bound $\|Pv-u\|\leq(\delta/\gamma)\sqrt{n-m}$ will hold.
@@
Consider the basis of eigenvectors of $L$. The Laplacian is symmetric and therefore diagonalizable since the graph is undirected. Therefore its eigenvalues form an orthogonal basis. Let $\{v_{1},\dotsc,v_{n}\}$ be a set of normal representatives, forming the spectral basis of $L$.
@@
By the assumption $\|\epsilon\|\leq\delta$ it follows $|\beta_{i}|\leq\delta$. Suppose that $H$ is the basis transformation matrix for the eigenbasis of $L$ (its rows are the $v_{i}$). Then $\|\beta\|^{2}=\langle\beta,\beta\rangle=\langle H\epsilon,H\epsilon\rangle=\langle\epsilon,H^{T}H\epsilon\rangle=\langle\epsilon,\epsilon\rangle=\|\epsilon\|^{2}$ since $\langle\cdot,\cdot\rangle$ is a norm and $H$ is orthonormal. Thus $\|\epsilon\|=\|\beta\|$ and therefore any bound on the norm of $\epsilon$ carries over to the norm of $\beta$. Since $\|\cdot\|$ denotes the euclidean norm, it follows $|\beta_{i}|\leq\|\beta\|$ and therefore that $|\beta_{i}|\leq\delta$ for $1\leq i\leq n$
Recall that $L$ is positive semidefinite, and therefore has only non negative eigenvalues. From the above result we have that $\beta_{i}=(\lambda-\lambda_{i})\alpha_{i}$ for $1\leq i\leq n$. Thus $|(\lambda-\lambda_{i})\alpha_{i}|\leq\delta$ for all $i$.
It follows that when $\delta$ is small, $(\lambda-\lambda_{i})$ or $\alpha_{i}$ is small. For a fixed $\beta_{i}$ value, the further away $\lambda_{i}$ is from $\lambda$, the smaller $\alpha_{i}$ must be. This means that eigenvectors of $L$ with eigenvalues close to $\lambda$ can have larger coefficients $\alpha_{i}$ and therefore contain the most information of $w$. Let $A=\{a_{1},\cdots,a_{m}\}$ be the set of coefficients for eigenvalues within $\gamma$ of $\lambda$. For all other eigenvalues, $\alpha_{i}$ is likely small. Therefore $u=\alpha_{a_{1}}v_{a_{1}}+\cdots+\alpha_{a_{m}}v_{a_{m}}$ is our approximation of $Pv$ in the spectral basis. This allows for the derivation of a formal bound.
@@
This characterizes an eigenvector whose eigenvalue is close to that of $Pv$ as being a good approximation of it. As $\|\epsilon\|$ approaches zero, $\|Pv-u\|$ does as well for any truncated approximation $u$. In particular, when $\delta$ is small, then with a $\gamma$ small enough to only encompass a single eigenvalue, the corresponding eigenvector approximates $Pv$ well. This shows a stability in the spectral properties of the almost equitable partition. We call these approximations of AEPs, quasi-equitable partitions ($\delta$-QEPs). In the specific, define a $\delta$-QEP or a QEP at level $\delta$, to be a partition $\pi$ such that $\|E\|\leq\delta$ where $E$ is the deviation matrix $PL^{\pi}-LP$. We show an example at $\delta=2.4$ in figure 5 A. Whereas Kato and IshiiKato and Ishii (2023) define their weakening in terms of a bound on the edge weight sum of vertices to partition cells, we define ours with the norm of a matrix capturing the deviation from the average, which allows it to maintain a relationship with the underlying spectral theory as we have demonstrated above.
@@ III.2 Weak Cluster Synchronization
Because our description of cluster synchronization is spectral instead of combinatorial, we can immediately apply the same methodology from AEPs to the structural eigenvectors of $\delta$-QEPs to describe synchronized clusters (figure 5). When moving from an AEP to a $\delta$-QEP, we can no longer construct the full eigenbasis of a corresponding quotient Laplacian. Non-structural eigenvectors have nonzero coefficients in the eigendecomposition of a natural frequency vector, $\vec{\omega}$, with constant values within clusters. Information is leaked into the rest of the eigenbasis, and it is precisely this leakage which results in differences in phase within clusters, as seen in figure 5 B. This is small if the $\delta$-QEP is reasonably close to an AEP. In this case the contribution of the non structural eigenvectors in the description of the natural frequencies results in a small but nonzero end state coefficient for them (figure 5 C.). Additionally, the structural eigenvectors are no longer completely constant within clusters. The non exact cluster synchronization arises from the difference in scaling moving from the decomposition of the natural frequency vector to the decomposition of the end state of the system.
Figure 5: A. The weighted network consists of $15$ vertices and a color coded $\delta$-QEP at level $\delta=2.4$ which it admits. This graph was generated by adding uniform noise to the toy example from figure 2. B. The phase plot indicates the phase locked cluster synchronization of the Kuramoto model placed on this graph with coupling $\sigma=1.2$ and identical natural frequencies within each cell (respectfully the pink, yellow, blue cells have frequencies $3,4,5$). This frequency distribution reinforces cluster synchronization by mostly exciting structural modes of the $\delta-QEP$. C.Time evolution of the coefficients in the projection onto the Laplacian eigenbasis. The cluster synchronized state is described almost entirely by the quotient graph sub-basis, $\alpha_{1}(t)$ and $\alpha_{2}(t)$, with the leakage of information to the rest of the eigenbasis indicated by the spread of non structural coefficients around zero in C., and slight spread of phases within clusters in B.
In section II.1.1 we discussed assortativity and spectrum placement with AEPs. The same argument using random walks can be extended to weighted graphs and $\delta$-QEPs.Saade, Krzakala, and Zdeborová (2014) The proximity of a $\delta$-QEP to a full AEP determines the bound of the variance in the intra cluster values of a structural eigenvector. With a reasonably small $\delta$, we can expect generically few zero crossings to occur within a cluster. Further, since proposition 8 found that the best truncated approximation of an AEP structural eigenvector in the spectrum of a close $\delta$-QEP uses the eigenvectors with the nearest eigenvalues, we can expect the transient clustering discussed in section II.4 to behave well in the generalization to $\delta$-QEPs, as the modes needed to describe clusters at a particular scale have close eigenvalues to one another and will consequently decay at a similar rate.
@@ IV.1 Clustered Dynamics at Separate Frequencies
An analytically tractable way to explore this through the coefficient equations is by considering the single unstable mode dynamics. Consider a coefficient $1\leq r_{1}\leq n-1$, and assume non zero modes $r\neq 0,r_{1}$ have settled into the asymptotic state predicted in the linearized case $\alpha_{r}^{\infty}$. We extend the derivation from KallionatisKalloniatis (2010) to weighted graphs. One can find parameter values of a regime change for the dynamical mode $\alpha_{r_{1}}$ in a second order approximation. Particularly we see the following, the derivation of which is in appendix C.
@@
Proposition 9.
Consider a single dynamical mode, $r_{1}$ as in the above derivation, corresponding to an AEP structural eigenvector without a repeated eigenvalue. Let $\vec{\omega}=av_{r_{1}}$ for $a\in\mathbb{R}$. Under these assumptions, the system experiences clustered behavior with at least two clusters at separate instantaneous frequencies.
Proof.
Let $\pi=\{V_{1},\dotsc,V_{k}\}$ be the AEP inducing the structural $v_{r_{1}}$. Since $\vec{\omega}=av_{r_{1}}$, it follows that $\alpha_{r}^{\infty}=0$ for all $r\neq r_{1},0$. Consequently $\vec{\theta}=\alpha_{r_{1}}(t)v_{r_{1}}$, which implies $\dot{\vec{\theta}}=\dot{\alpha}_{r_{1}}(t)v_{r_{1}}$.
It follows that $\dot{\theta}_{i}(t)=\dot{\alpha}_{r_{1}}(t)(v_{r_{1}})_{i}$. By the assumption on $v_{r_{1}}$, it follows that $(v_{r_{1}})_{i}=(v_{r_{1}})_{j}$ for all $i,j\in V_{p}$, where $1\leq p\leq k$. Thus all vertices within a cluster experience the same angular velocity. Since $r_{1}\neq 0$, there exist two partition cells such that their corresponding values in the vector $v_{r_{1}}$ differ. Thus for all $i,j\in V_{p}$ with $1\leq p\leq k$, we have $\dot{\theta}_{i}(t)=\dot{\theta}_{j}(t)$, and for some pair of $1\leq p,q\leq k$ we have $\dot{\theta}_{i}(t)\neq\dot{\theta}_{j}(t)$ for any $i\in V_{p}$ and $j\in V_{q}$.
Dividing the angular velocities by $2\pi$, we obtain the desired result for the angular frequency. ∎
@@
This shows that the lower spatial frequencies can support dynamical modes better. In the context of our analysis of structural eigenvectors, this suggests that the eigenvectors of assortative AEPs or $\delta$-QEPs allow for unstable mode behavior more easily. Thus clusters with strong intra cluster connectivity and weak inter cluster connectivity should be able to maintain separate cluster frequencies if the disparity in connectivity is high enough and the natural frequencies align well enough with one of the structural eigenvectors. This intuition also suggests that an analogous property should hold for multiple unstable modes, if they all share reasonably high alignment with the natural frequency vector.
@@
Figure 6: We dive deeper into clustered dynamics of the sample graph in Fig.2. A. The discriminants $\Delta_{r}$ for all eigenmodes $r$ as functions of the coupling $\sigma$, with fixed natural frequencies, constant within clusters ($\vec{\omega}$ is a multiple of the eigenvector $v_{1}$). B. The phase plots show the system experiencing multi-frequency clustered behavior with a choice of $\sigma=0.2$, inside the negative range of $\Delta_{1}$. Finally, C. depicts the coefficients of the phase decomposition, indicating a single dynamical mode, $\alpha_{1}(t)$. Note the periodic shifts in $\alpha_{1}(t)$ along with the activation and decay of a second mode ($\alpha_{2}(t)$). The periodic jumps in $\alpha_{2}(t)$ are reflected in the periodic changes in the clustered dynamics in B, but we note that while this is a periodically recurrent change, the clustered oscillators do experience a different number of oscillations from one another. In this case, the yellow cluster never completes a full oscillation, and the other two clusters are in counter rotation to one another.
This reasoning does not extend easily to $\delta$-QEPs. Certainly the machinery works for any eigenvectors of the graph Laplacian, but the information leak of the structural eigenvectors for a $\delta$-QEP complicates the description of clustered behavior. Given a system with constant natural frequencies within clusters, the structural eigenvectors of an associated $\delta$-QEP generically won’t be able to fully describe the clustered dynamics. In the frequency locked synchronization regime, this resulted in small activation from other modes. When considering a single unstable mode, that information from non structural modes can allow them to also become unstable. This behavior distances the dynamics from the above analysis, and makes the results less interpretable. However, the clustered behavior is still present, and a more catered choice of natural frequencies should allow for the same behavior in isolated coefficients.